OpenStudy (anonymous):

Can someone please check my answer? thank you!! 9a=11 a= 1.22

OpenStudy (ghostgate):

That should be correct. When it comes to solving the unknown variable you have to do the opposite of what's going on. Since "9a" is multiplication you divide "9" by both sides. 9a/9 = 11/9 1a = 1.22_ -or- a = 1.22 I hope that helps. ;) {-Ghostgate-}

OpenStudy (ghostgate):

Medal? I guess not.